OpenStudy (anonymous):

Evaluate the given limit lim x->0 of 1-cos(5x)/sin(9x)

hartnn (hartnn):

multiply and divide by (1+cos5x)

OpenStudy (anonymous):

now i have 1-cos(5x)^2 / sin(9x) x (1+cos5x)

OpenStudy (anonymous):

i am confused on how i would further simplify that.

hartnn (hartnn):

1-cos^2 x = ?

OpenStudy (anonymous):

sin^2x

hartnn (hartnn):

so 1-cos(5x)^2 = sin^2 5x, isn't it ? now multiply and divide by x

OpenStudy (anonymous):

5 lim x->0 of \[\sin ^{2}x/\sin9x \times1-\cos(5x)\]

OpenStudy (anonymous):

correct?

hartnn (hartnn):

wouldn't it be sin^2 (5x) ?? and (1+cos 5x) ?

OpenStudy (anonymous):

that is what i had, then i multiplied by 5x/5x so sin^2(5x) turns into one and im not sure what to do with the denominator

hartnn (hartnn):

sin^2 5x doesn't turn into 1 \(\large \frac{sin^25x}{sin 9x(1+cos 5x)}= \frac{5(sin^25x)/5x}{9(sin 9x)/9x}*\frac{1}{(1+cos 5x)}\) got this step?

OpenStudy (anonymous):

then i multiply by the conjugate?

hartnn (hartnn):

no.

OpenStudy (anonymous):

i got it. i need to multiply each sin by the number located by it to make each one equal one since lim x->0 sinx/x=1

hartnn (hartnn):

yes.

OpenStudy (anonymous):

5sin5x/(9sin9x)(1-cos5x) = 1*1/1+cos5x

OpenStudy (anonymous):

=1/1+cos5x

OpenStudy (anonymous):

now i don't know where to go from here.

hartnn (hartnn):

its sin^2 (5x) not sin 5x. \(\large\frac{5(sin^25x)/5x}{9(sin 9x)/9x}*\frac{1}{(1+cos 5x)}=\frac{5sin 5x (1)}{9(1)}\frac{1}{1+cos 5x}\) did u get this ? now put x=0, directly

OpenStudy (anonymous):

5/9 * 1/1+1 =5/9 * 1/2 = 5/18 ?

hartnn (hartnn):

when u put x= 0 what is sin 0 ??

OpenStudy (anonymous):

0

hartnn (hartnn):

so? wouldn't the answer be 0 then ?

OpenStudy (anonymous):

correct because if you have 0 in the numerator than the whole equation turns to zero, correct?

hartnn (hartnn):

yup

OpenStudy (anonymous):

you end up multiplying the whole equation by zero, meaning the answer is 0?

hartnn (hartnn):

yes!

OpenStudy (anonymous):

thank-you very much! greatly appreciated, you were a great help.

hartnn (hartnn):

welcome ^_^ did u understand all steps? if doubts anywhere, just ask.

OpenStudy (anonymous):

i will send you a message if i need any other help but it is a lot clearer to me now. thanks a lot.

hartnn (hartnn):

ok :)